OpenStudy (samsan9):

log(10^3x)=? A.) x B.) 10^3 C.) 3x D.) 30

OpenStudy (tanya123):

1000x?

OpenStudy (samsan9):

this question makes little sense to me this is all it says

hartnn (hartnn):

\(\Large \log (10^{3x})\) use the property, \(\Large \log a^b = b \log a\)

hartnn (hartnn):

and since base of log is 10, we can write \(\log 10=1\)

OpenStudy (samsan9):

but what does this question mean i don;t get it there is nothing here asking me to find anything except the ?

hartnn (hartnn):

the question here is asking you the equivalent expression for log(10^3x) or in other words, it asks you to simplify log(10^3x)

hartnn (hartnn):

now we can simplify \(\log a^b\) as \(\Large b \log a\) Note that the exponent of 'a' now became co-efficient of log a this is one beautiful property of log which you need to apply for this problem :)

OpenStudy (samsan9):